# LATTICES OF FLATS FOR SYMPLECTIC MATROIDS

### v5

@@ 3.7. Theorem:
If $I(\mathcal{L},P)$ is the family of independent sets of an ordinary matroid then $\mathcal{L}$ can be uniquely extended to a geometric lattice $\mathcal{P}$, denoted the induced geometric lattice of $\mathcal{L}$, with $I\left(\mathcal{L},P\right)=I\left(\mathcal{P},P\right)$. In this case the restriction $f\mid_{\mathcal{L}}$ of the rank function $f$ of $\mathcal{P}$ is the original rank function on $\mathcal{L}$.
@@
We know that there exist a unique ordinary matroid $M_{\mathcal{L}}$ on the ground set $A(L)$ corresponding to the family of independent sets $I\left(\mathcal{L},P\right)$ of $\mathcal{L}$. We identify each element $X\in\mathcal{L}$ with the set $D=\left\{a\in A(\mathcal{L})\mid a\leq X\right\}$. To see $D$ is a flat of $M_{\mathcal{L}}$ observe that for every atom $a\notin D$ we have $X\lneq\vee(D\cup\left\{a\right\})$ and so
$rank(X)\lneq rank(D\cup\left\{a\right\})$
As the rank function of the matroid agrees with the rank function of $\mathcal{L}$ we must then have the lattice of flats of $M_{\mathcal{L}}$ be the desired geometric lattice.
@@ 3.8. Remark:
It is not always true that $I\left(\mathcal{L},P\right)$ is a family of independent sets of an ordinary matroid. A wide class of examples can be found in the study of the adjoint of a matroid. The first example I could find is by A. C. Cheung and can be observed in [10]. Another known example is the dual lattice of the Vamos matroid.
In addition to the previous embedding theorem, we give a characterization of geometric lattices using independent sets. We call an independent set $I$ geometric if $\left|I\right|=rank(I)$.
@@ 3.9. Theorem:
This property is known for the geometric lattice corresponding to the ordinary matroid $I\left(\mathcal{L},P\right)$. Let $X,Y\in\mathcal{L}$ and $I$ be an independent set that spans $X\wedge Y$. We can extend $I$ to independent sets $I\cup J$ that span $X$ and $I\cup J^{\prime}$ that span $Y$. Note also that $I\cup J\cup J^{\prime}$ spans $X\vee Y$ and therefore contains an independent set $K$ that spans $X\vee Y$. This proves that $\mathcal{L}$ is sub-modular as:
@@
In figure [3.4] we can see an example of the induced geometric lattice of the dual lattice of the partition lattice on the set of $4$ elements. The element $X$ added to the lattice reduces the rank of the size two subsets of $\left\{124/3,13/24,123/4\right\}$, which are the only independent sets with different size and rank.
@@ 4.1. Remark:
We move on to showing the correspondence between $C_{n}$ lattices and ranked symplectic matroids, starting with some definitions.
@@ 4.3. Definition:
    If $I_{1}$ and $I_{2}$ are members of $\mathcal{I}$ such that $\left|I_{1}\right|\lneq\left|I_{2}\right|$, then either there exists $a\in I_{2}\setminus I_{1}$ such that $\left\{a\right\}\cup I_{1}\in\mathcal{I}$ or there exists $a\notin I_{1}\cup I_{2}$ such that both $\left\{a\right\}\cup I_{1}\in\mathcal{I}$ and $\left\{a^{*}\right\}\cup I_{1}\setminus{I_{2}}^{*}\in\mathcal{I}$.
In this section we use a slightly different notion of $NBB$ and independent sets defined in section 2. The reason is that we want independent sets to be subsets of the ground set $J$ and not of the set of atoms $A(\mathcal{L})$. For a subset $X\subseteq J$ we define $A(X)\subseteq A(\mathcal{L})$ to be the set of atoms that contain the elements of $X$, we say $X$ is disjoint in $\mathcal{L}$ if $A(X)\neq A(X\setminus\left\{a\right\})$ for all $a\in X$. Notice that a linear order on $J$ induces a linear orders on the set of atoms of a $C_{n}$ lattice, giving the desired connection between the definitions. All of the theorems in the previous section hold with a small change in formulation using the new definitions.
@@ 4.4. Definition:
Let $(\mathcal{L},\omega)$ be a bounded lattice of subsets on a finite ground set $E$ with a partial order $\omega$ on $E$. A nonempty disjoint in $\mathcal{L}$ set $D\subseteq E$ is bounded below or $BB$ if there exist $a\in E$ such that $a$ is a strict lower bound for all $d\in D$ in the order $\omega$ and $a\in\vee D$.
@@
We define the family of independent sets $I\left(\mathcal{L},P\right)$ of a bounded lattice $\mathcal{L}$ on the ground set $E$ with $P$ the set of all linear orders on $E$.
$I\left(\mathcal{L},P\right)=\left\{A\in P\left(J\right)\mid\exists\omega\in P\text{ s.t. }A\text{ is an }NBB\text{ set in }\left(\mathcal{L},\omega\right)\right\}$
For a $C_{n}$ lattice $\mathcal{L}$, for which the ground set is $J$, we denote by $I^{ad}\left(\mathcal{L},P\right)$ the family of admissible independent sets of $\mathcal{L}$ and by $I^{\overline{ad}}\left(\mathcal{L},P\right)$ the family of admissible sets of $\mathcal{L}$ with admissible join. We are ready to state the main theorem of this section:
@@ 4.5. Theorem:
If $\mathcal{L}$ is a $C_{n}$ lattice then $I^{ad}\left(\mathcal{L},P\right)$ is a family of independent sets of a symplectic matroid $\mathcal{B}$.
In order to prove this theorem we need to work with two types of ordinary matroids induced by $\mathcal{L}$. We start with the first type of ordinary matroid, corresponding to the family of independent sets of the lattice $\mathcal{L}\cap T=\left\{X\cap T\mid X\in\mathcal{L}\right\}$ for a transversal $T$. We introduce the following lemma:
@@ 4.6. Lemma:
For any $A\in P^{ad}(J)$, we have $I\left(\mathcal{L},P\right)\cap A=I\left(\mathcal{L}\cap A,P\mid_{A}\right)$.
@@
For the first inclusion it is enough to prove that if $D$ is a minimal $BB$ set in $\mathcal{L}\cap A$ with respect to a partial order $\omega$ on $A$ then $D$ is a $BB$ in $\mathcal{L}$ with respect to any extension of $\omega$ to $J$. This is the case as the lower bound for $D$ in $\mathcal{L}\cap A$ will also be a lower bound in $\mathcal{L}$. For the second inclusion let $I$ be $NBB$ in $\mathcal{L}\cap A$ with respect to an order $\omega$, we have that with respect to a partial order starting with $\omega$, $I$ is $NBB$ in $\mathcal{L}$.
@@ 4.7. Lemma:
Let $I\in I^{ad}\left(\mathcal{L},P\right)$ and $a\in I$ then exactly one of $I\cup\left\{a^{*}\right\}\in I\left(\mathcal{L},P\right)$ or $rank(I)=rank(\mathcal{L})$ holds.
@@
Assume $I\cup\left\{a^{*}\right\}\notin I\left(\mathcal{L},P\right)$ then by definition [4.4] we have $a^{*}\in\vee I$. Therefore, $\vee I$ is inadmissible and we get from definition [2.1] that $rank(I)=rank(\mathcal{L})$.
We can conclude that $I\cup\left\{a,a^{*}\right\}\in I\left(\mathcal{L},P\right)$ for $I\in I^{ad}\left(\mathcal{L},P\right)$ iff $I\cup\left\{a\right\},I\cup\left\{a^{*}\right\}\in I^{\overline{ad}}\left(\mathcal{L},P\right)$. On the other hand, if $a,a^{*},b,b^{*}\in B\subseteq A(\mathcal{L})$ then $a\in\vee B\setminus\left\{a\right\}$ and $b\in\vee B\setminus\left\{b\right\}$, therefore $B$ is not an independent set. We have thus proven the following theorem:
@@ 4.8. Theorem:
$I\left(\mathcal{L},P\right)=I^{ad}\left(\mathcal{L},P\right)\sqcup\left\{I\cup\left\{a,a^{*}\right\}\mid I\cup\left\{a\right\},I\cup\left\{a^{*}\right\}\in I^{\overline{ad}}\left(\mathcal{L},P\right)\right\}$
The second type of ordinary matroid that we introduce is the one induced by the original family of independent sets of $\mathcal{L}$, including the inadmissible independent sets. We use Theorem [3.7] to induce a unique geometric lattice $\mathcal{P}$ on the same ground set with the same family of independent sets such that $\mathcal{L}\subseteq\mathcal{P}$ denotes the geometric lattice induced by $\mathcal{L}$.
@@ 4.9. Theorem:
If $\mathcal{L}$ is a $C_{n}$ lattice, then $I\left(\mathcal{L},P\right)$ is a family of independent sets of an ordinary matroid.
@@
We prove that $I\left(\mathcal{L},P\right)$ has the augmentation property for independent sets, stating that for a smaller independent set $I_{1}$ there exists $a\in I_{2}\setminus I_{1}$ such that $I_{1}\cup\left\{a\right\}$ is independent. Let $I_{1},I_{2}\in I\left(\mathcal{L},P\right)$ be such that $\left|I_{1}\right|\lneq\left|I_{2}\right|$. Notice that if $rank(I_{1})\neq rank(\mathcal{L})$, then $rank(I_{1})=\left|I_{1}\right|\lneq rank(I_{2})$, and we can therefore assume $rank(I_{1})=rank(\mathcal{L})$.
Let $a\in I_{1}$ be an atom such that $rank(I_{1}\setminus\left\{a\right\})\lneq rank(I_{1})$ is guaranteed by lemma [3.3]. It is enough to find an atom $b\in I_{2}$ for which the following strict inequality holds:
$rank(I_{1}\setminus\left\{a\right\})\lneq rank((I_{1}\setminus\left\{a\right\})\cup\left\{b\right\})\lneq rank(\mathcal{L})$
Following lemma [4.7], the inequality above holds for $b\in I_{2}$ iff $b,b^{*}\notin\vee(I_{1}\setminus\left\{a\right\})$. As we have subtracted an element from $I_{1}$ we can do the same for $I_{2}$ while keeping it larger, so assume $rank(I_{2})\lneq rank(\mathcal{L})$. Therefore, applying lemma [4.7] we have $(I_{2}\setminus\left\{b\right\})\cup\left\{b^{*}\right\}\in I^{ad}\left(\mathcal{L},P\right)$ and $rank(I_{2}\setminus\left\{b\right\})\cup\left\{b^{*}\right\}))\lneq rank(\mathcal{L})$ for every $b\in I_{2}$. If there exist no $b\in I_{2}$ such that $b,b^{*}\notin\vee(I_{1}\setminus\left\{a\right\})$ then by replacing all $b\in I_{2}$ with $b^{*}\in\vee(I_{1}\setminus\left\{a\right\})$ we obtain an admissible independent set $I_{3}$ such that $\vee I_{3}\subseteq\vee I_{1}\subsetneq J$ but $\left|I_{1}\right|\lneq\left|I_{2}\right|=\left|I_{3}\right|$, contradicting Lemma [3.4].

### v6

@@ 3.7. Theorem:
If $I(\mathcal{L})$ is the family of independent sets of an ordinary matroid, then $\mathcal{L}$ can be uniquely extended to a geometric lattice $\mathcal{P}$, denoted the induced geometric lattice of $\mathcal{L}$, with $I\left(\mathcal{L}\right)=I\left(\mathcal{P}\right)$. In this case, the restriction $f\mid_{\mathcal{L}}$ of the rank function $f$ of $\mathcal{P}$ is the original rank function on $\mathcal{L}$.
@@
We know that there exists a unique ordinary matroid $M$ on the ground set $A(\mathcal{L})$ that corresponds to the family of independent sets $I\left(\mathcal{L}\right)$ of $\mathcal{L}$. We identify each element $X\in\mathcal{L}$ with the set $D=\left\{a\in A(\mathcal{L})\mid a\leq X\right\}$. Using Lemma 3.6 we know that $rank_{\mathcal{L}}(X)$ is the cardinality of a maximal independent set contained in $X$. The same is true for $rank_{\mathcal{M}}(D)$ by definition. As $I\left(\mathcal{L}\right)=I\left(\mathcal{P}\right)$, we must have $rank_{\mathcal{L}}(X)=rank_{\mathcal{M}}(D)$.
To see $D$ is a flat of $M_{\mathcal{L}}$ observe that for every atom $a\notin D$ we have $X\lneq\vee(D\cup\left\{a\right\})$ and so:
$rank_{\mathcal{L}}(D)=rank_{\mathcal{L}}(X)\lneq rank_{\mathcal{L}}(D\cup\left\{a\right\})=rank_{\mathcal{M}}(D\cup\left\{a\right\})$
@@ 3.8. Remark:
It is not always true that $I\left(\mathcal{L}\right)$ is a family of independent sets of an ordinary matroid. A wide class of examples can be found in the study of the adjoint of a matroid. The first example I could find is by A. C. Cheung and can be observed in [10]. Another known example is the dual lattice of the Vamos matroid.
In addition to the previous embedding Theorem, we can give a characterization of geometric lattices using independent sets. We call an independent set $I$ geometric if $\left|I\right|=rank(I)$.
@@ 3.9. Theorem:
If $\mathcal{L}$ is geometric, then $I\left(\mathcal{L}\right)$ is the family of independent sets of a matroid, and so every $I\in I\left(\mathcal{L}\right)$ is geometric. On the other hand, we only need to prove that $\mathcal{L}$ is submodular. Let $X,Y\in\mathcal{L}$ and $I$ be an independent set that spans $X\wedge Y$. Using lemma 3.6 we can extend $I$ to independent sets $I\cup J$ that span $X$ and $I\cup J^{\prime}$ that span $Y$. Note also that $I\cup J\cup J^{\prime}$ spans $X\vee Y$ and therefore contains an independent set $K$ that spans $X\vee Y$. Using the fact that every independent set is geometric, we obtain the following.
  1. (1)
    $rank(X\wedge Y)=\left|I\right|$
  2. (2)
    $rank(X\vee Y)=\left|K\right|\leq\left|I\right|+\left|J\right|+\left|J^{\prime}\right|$
This is enough to prove submodularity as follows:
@@
In Figure 3.10 we can see an example of the induced geometric lattice. The partition lattice on the set of $4$ elements is a geometric lattice. Although its dual lattice $(a)$, is not geometric, it is finite, atomic, and graded. Moreover, its family of independent sets constitutes a matroid and therefore can be extended to its induced geometric lattice $(b)$. The element $X$ added to the lattice reduces the rank of the size two subsets of $\left\{14/23,13/24,12/34\right\}$, which are the only non geometric independent sets.
1/2/3/41/2/3/41/23/41/23/414/2/314/2/31/24/31/24/313/2/413/2/412/3/412/3/41/2/341/2/3414/2314/231/2341/234124/3124/313/2413/24123/4123/4134/2134/212/3412/3412341234(a)(a)
1/2/3/41/2/3/41/23/41/23/414/2/314/2/31/24/31/24/313/2/413/2/412/3/412/3/41/2/341/2/34XX14/2314/231/2341/234124/3124/313/2413/24123/4123/4134/2134/212/3412/3412341234(b)(b)
3.10. Figure:
@@ 4.1. Remark:
We move on to showing the correspondence between $C_{n}$ lattices and ranked symplectic matroids, starting with some definitions following [3].
@@ 4.3. Definition:
    If $I_{1}$ and $I_{2}$ are members of $\mathcal{I}$ such that $\left|I_{1}\right|\lneq\left|I_{2}\right|$, then either there exists $a\in I_{2}\setminus I_{1}$ such that $\left\{a\right\}\cup I_{1}\in\mathcal{I}$ or there exists $a\notin I_{1}\cup I_{2}$ such that both $\left\{a\right\}\cup I_{1}\in\mathcal{I}$ and $(\left\{a^{*}\right\}\cup I_{1})\setminus{I_{2}}^{*}\in\mathcal{I}$.
In this section, we use a slightly different notion of $NBB$ and independent sets defined in section 2. The reason is that we want the independent sets to be subsets of the ground set $J$ and not of the atom set $A(\mathcal{L})$. For a subset $X\subseteq J$ we define $A(X)\subseteq A(\mathcal{L})$ to be $\left\{B\in A(\mathcal{L})\mid B\cap X\neq\emptyset\right\}$, that is, the set of atoms that contain the elements of $X$. We say $X$ is disjoint in $\mathcal{L}$ if the intersection of $X$ with any atom of $\mathcal{L}$ is empty or a singleton.
Given a linear order $\omega$ on $J$ we also denote by $\omega$ the following linear orders on the set of atoms of a $C_{n}$ lattice $\mathcal{L}$. $B\lneq_{\omega}C$ for $B,C\in A(\mathcal{L})$ if $min_{\omega}B\leq min_{\omega}C$. These changes will give us the desired connection between the definitions. We now reformulate the results of Section 3 in this setting.
@@ 4.4. Definition:
Let $(\mathcal{L},\omega)$ be a finite lattice of subsets on the ground set $E$ with a partial order $\omega$ on $E$. A nonempty disjoint in $\mathcal{L}$ set $D\subseteq E$ is bounded below or $BB$ if there exist $a\in E$ such that $a$ is a strict lower bound for all $d\in D$ in the order $\omega$ and $a\in\vee D$.
@@
We define the family of independent sets $I\left(\mathcal{L}\right)$ of a finite lattice $\mathcal{L}$ on the ground set $E$ with $T$ the set of all linear orders on $E$.
[…]
For a $C_{n}$ lattice $\mathcal{L}$, for which the ground set is $J$, we denote by $I^{ad}\left(\mathcal{L}\right)$ the family of admissible independent sets of $\mathcal{L}$ and by $I^{\overline{ad}}\left(\mathcal{L}\right)$ the family of admissible sets of $\mathcal{L}$ with admissible join. We are ready to state the main Theorem of this section:
@@ 4.5. Theorem:
If $\mathcal{L}$ is a $C_{n}$ lattice, then $I^{ad}\left(\mathcal{L}\right)$ is the family of independent sets of a symplectic matroid $\mathcal{B}$.
Example:
The Family of independent sets $I(\mathcal{L}(U^{*}_{k,n}))$ for the $C_{n}$ lattice in Example 2 is the family of all admissible sets of size no grater than $t$.
To prove this Theorem, we will need to work with two types of ordinary matroids induced by $\mathcal{L}$. We start with the first type of ordinary matroid, corresponding to the family of independent sets of the lattice $\mathcal{L}\cap A=\left\{X\cap A\mid X\in\mathcal{L}\right\}$ for an admissible $A\in 2^{J}$. We introduce the following Lemma:
@@ 4.6. Lemma:
For any admissible $A\in 2^{J}$, we have $I\left(\mathcal{L}\right)\cap A=I\left(\mathcal{L}\cap A\right)$.
@@
For the inclusion $\subseteq$, it suffices to prove that if $D$ is a minimal $BB$ set in $\mathcal{L}\cap A$ with respect to a partial order $\omega$ on $A$, then $D$ is a $BB$ in $\mathcal{L}$ with respect to any extension of $\omega$ to $J$. This is the case, as the lower bound for $D$ in $\mathcal{L}\cap A$ will also be a lower bound in $\mathcal{L}$. For the inclusion $\supseteq$, let $I$ be $NBB$ in $\mathcal{L}\cap A$ with respect to an order $\omega$, we have that with respect to a partial order starting with $\omega$, $I$ is $NBB$ in $\mathcal{L}$.
4.7. Corollary:
For any admissible $A\in 2^{J}$, we have $I\left(\mathcal{L}\right)\cap A$ the family of independent sets of a matroid.
@@ 4.8. Lemma:
Let $I\in I\left(\mathcal{L}\right)$ and $a\in I$ then exactly one of $I\cup\left\{a^{*}\right\}\in I\left(\mathcal{L}\right)$ or $rank(I)=rank(\mathcal{L})$ holds.
@@
Assume $I\cup\left\{a^{*}\right\}\notin I\left(\mathcal{L}\right)$, then by Lemma 3.5 we have $a^{*}\in\vee I$. Therefore, $\vee I$ is inadmissible and we get from Definition 2.3 that $rank(I)=rank(\mathcal{L})$.
@@ 4.9. Theorem:
$I\left(\mathcal{L}\right)=I^{ad}\left(\mathcal{L}\right)\sqcup\left\{I\cup\left\{a,a^{*}\right\}\mid I\cup\left\{a\right\},I\cup\left\{a^{*}\right\}\in I^{\overline{ad}}\left(\mathcal{L}\right)\right\}$
Proof:
We first observe that an inadmissible independent set of $\mathcal{L}$ contains exactly one inadmissible pair. If $a,a^{*},b,b^{*}\in B\subseteq A(\mathcal{L})$, then $a\in\vee B\setminus\left\{a\right\}$ and $b\in\vee B\setminus\left\{b\right\}$, therefore $B$ is not an independent set.
Next, let $I\cup\left\{a,a^{*}\right\}\in I\left(\mathcal{L}\right)$ with $I\in I^{ad}\left(\mathcal{L}\right)$. Using Lemma 4.8, we have $rank(I\cup\left\{a\right\}),rank(I\cup\left\{a^{*}\right\})\neq rank(\mathcal{L})$ and so $I\cup\left\{a\right\},I\cup\left\{a^{*}\right\}\in I^{\overline{ad}}$. On the other hand, if $I\cup\left\{a\right\},I\cup\left\{a^{*}\right\}\in I^{\overline{ad}}$, then again by Lemma [4.8] we get $I\cup\left\{a,a^{*}\right\}\in I\left(\mathcal{L}\right)$.
The second type of ordinary matroid that we introduce is the one induced by the original family of independent sets of $\mathcal{L}$, including the inadmissible independent sets. Reformulating Theorem 3.7, we will induce a unique geometric lattice $\mathcal{P}$ on $J$ with the same family of independent sets. For that purpose, we first introduce another useful technical Lemma.
4.10. Lemma:
Let $\mathcal{L}$ be a $C_{n}$ lattice, if $I\in I^{\overline{ad}}(\mathcal{L})$, then $rank(I)=\left|I\right|$. Moreover, $(I\setminus\left\{a\right\})\cup\left\{a^{*}\right\}\in I^{\overline{ad}}(\mathcal{L})$ for every $a\in I$.
Proof:
Let $I\in I^{\overline{ad}}(\mathcal{L})$, if $rank(I)\neq\left|I\right|$ then there exist $I^{\prime}\subsetneq I$ and $a\in I\setminus I^{\prime}$ such that $rank(I^{\prime})=\left|I^{\prime}\right|$ but $rank(I^{\prime}\cup\left\{a\right\})\neq\left|I^{\prime}\cup\left\{a\right\}\right|=rank(I^{\prime})+1$. Applying Lemma 3.4 we must have $rank(I^{\prime}\cup\left\{a\right\})\gneq rank(I^{\prime})+1$. As $I^{\prime}\cup\left\{a\right\}\subseteq I$, we have $I^{\prime}\cup\left\{a\right\}\in I^{\overline{ad}}(\mathcal{L})$. Therefore, $a,a^{*}\notin\vee I^{\prime}$ and by property $4$ of Definition 2.3 we obtain an element $A\in\mathcal{L}$ covering $\vee I^{\prime}$ and containing $a$. We must have $\vee(I^{\prime}\cup\left\{a\right\})=A$ contradicting $rank(I^{\prime}\cup\left\{a\right\})\gneq rank(I^{\prime})+1$.
For the second part, let $a\in I\in I^{\overline{ad}}(\mathcal{L})$, we again use property $4$ of Definition 2.3 to obtain an element $A\in\mathcal{L}$ covering $I\setminus\left\{a\right\}$ and containing $a^{*}$ with $A\cap(\vee I)=\vee(I\setminus\left\{a\right\})$. We therefore have $\vee((I\setminus\left\{a\right\})\cup\left\{a^{*}\right\})=A\neq J$ and so $(I\setminus\left\{a\right\})\cup\left\{a^{*}\right\}\in I^{\overline{ad}}(\mathcal{L})$.
The first step to finding the induced geometric lattice $\mathcal{P}$ is to check the hypothesis of Theorem 3.7, as follows:
@@ 4.11. Theorem:
If $\mathcal{L}$ is a $C_{n}$ lattice, then $I\left(\mathcal{L}\right)$ is a family of independent sets of an ordinary matroid.
@@
We prove that $I\left(\mathcal{L},P\right)$ admits the augmentation property for independent sets, stating that for two independent sets $I_{1},I_{2}$ with $\left|I_{1}\right|\lneq\left|I_{2}\right|$ there exists $a\in I_{2}\setminus I_{1}$ such that $I_{1}\cup\left\{a\right\}$ is independent.
Let $I_{1},I_{2}\in I\left(\mathcal{L}\right)$ be such that $\left|I_{1}\right|\lneq\left|I_{2}\right|$. Observe that if $I_{1}\in I^{\overline{ad}}(\mathcal{L})$, then by Lemmas 3.4 and 4.10 we have that $rank(I_{1})=\left|I_{1}\right|\lneq\left|I_{2}\right|\leq rank(I_{2})$. Therefore, there exist $a\in I_{2}$ such that $a\notin\vee I_{1}$ and by Lemma 3.5 $I_{1}\cup\left\{a\right\}$ is independent.
If $I_{1}\notin I^{\overline{ad}}(\mathcal{L})$, let $b\in I_{1}$ be such that $rank(I_{1}\setminus\left\{b\right\})\lneq rank(I_{1})$ using Lemma 3.5. We again use Lemma 3.5 to assume that $I_{2}\in I^{\overline{ad}}(\mathcal{L})$, letting it be smaller but still having $\left|I_{1}\setminus\left\{b\right\}\right|\lneq\left|I_{2}\right|$. We define:
$I_{2}^{\prime}=\left\{a\mid(a\in\vee(I_{1}\setminus\left\{b\right\}))\,and\,(\left\{a,a^{*}\right\}\cap I_{2}\neq\emptyset)\right\}$
By Theorem 4.11 we have $I_{2}^{\prime}\in I^{\overline{ad}}(\mathcal{L})$ as it is obtained by a sequence of replacing an element with its star. We have the following:
$\left|I_{2}^{\prime}\right|=rank(I_{2}^{\prime})\leq rank(I_{1}\setminus\left\{b\right\})=\left|I_{1}\setminus\left\{b\right\}\right|\lneq\left|I_{2}\right|$
Therefore, there exists $a\in I_{2}$ such that $a,a^{*}\notin\vee(I_{1}\setminus\left\{b\right\})$. As $rank(I_{1}\setminus\left\{b\right\})\lneq rank(I_{2})\lneq rank(\mathcal{L})$ and $\vee((I_{1}\setminus\left\{b\right\})\cup\left\{a\right\})$ covers $\vee(I_{1}\setminus\left\{b\right\})$ we have $(I_{1}\setminus\left\{b\right\})\cup\left\{a\right\}\in I^{\overline{ad}}(\mathcal{L})$. Finally:
$rank(((I_{1}\setminus\left\{b\right\})\cup\left\{a\right\}))\lneq rank(\mathcal{L})=rank(I_{1})\leq rank(I_{1}\cup\left\{a\right\})$
then by Lemma 3.5 $I_{1}\cup\left\{a\right\}\in I(\mathcal{L})$.
Proof of Theorem 4.5:
By definition, $NBB$ is a hereditary property, so $I\left(\mathcal{L}\right)$ is a subset closed family.
  1. (1)
    Given a transversal $T\subset J$ we have by Lemma 4.6 that $I\left(\mathcal{L}\right)\cap T=I\left(\mathcal{L}\cap T\right)$. By Theorem 3.7 we then find that $I\left(\mathcal{L},P\right)\cap T$ is the family of independent sets of an ordinary matroid on the ground set $T$.
  2. (2)
    Let $I_{1},I_{2}\in I^{ad}\left(\mathcal{L}\right)$ be such that $\left|I_{1}\right|\lneq\left|I_{2}\right|$. Assume that for every $a\in I_{2}\setminus I_{1}$ such that $I_{1}\cup\left\{a\right\}\in I\left(\mathcal{L}\right)$ has $a^{*}\in I_{1}$. By Theorem 4.11, there exists such an $a$, and so by Lemma 4.8 we have $I_{1}\in I^{\overline{ad}}\left(\mathcal{L}\right)$. We now observe that $\left|I_{1}\setminus I_{2}^{*}\right|\lneq\left|I_{2}\setminus I_{1}^{*}\right|$ for the smaller independent sets $I_{1}\setminus I_{2}^{*},I_{2}\setminus I_{1}^{*}$. By Theorem 4.11 we have some $b\in I_{2}\setminus I_{1}^{*}$, $b\notin I_{1}\setminus I_{2}^{*}$ such that $(I_{1}\setminus I_{2}^{*})\cup\left\{b\right\}\in I\left(\mathcal{L}\right)$. We actually have $(I_{1}\setminus I_{2}^{*})\cup\left\{b\right\}\in I^{ad}\left(\mathcal{L}\right)$ because we have removed all inadmissible pairs in $I_{1}\cup I_{2}$. We have found $b\in I_{2}\setminus I_{1}$ with $b^{*}\notin I_{1}\cup I_{2}$. By our initial assumption $I_{1}\cup\left\{b\right\}\notin I\left(\mathcal{L}\right)$, and as $I_{1}\in I^{\overline{ad}}\left(\mathcal{L}\right)$ we have $I_{1}\cup\left\{b^{*}\right\}\in I\left(\mathcal{L}\right)$ by Lemma 3.5. As $b\notin I_{1}$, we get $I_{1}\cup\left\{b^{*}\right\}\in I^{ad}\left(\mathcal{L}\right)$ satisfying the second part of condition (2) of Definition 4.3.
After showing that every $C_{n}$ lattice corresponds to a symplectic matroid, we continue by identifying the symplectic matroids corresponding to $C_{n}$ lattices.